If the NPV (Net Present Value) of a shop is R195 000 and the profitability index is 1,24375, the initial investment in the shop is


[1] R86 908. [2] R800 000. [3] R195 000. [4] R156 784. [5] none of the above


1
Expert's answer
2021-07-20T16:05:35-0400

"PI=(NPV+" initial investment)"\div" initial investment

"1.24375=(195 000+Initial \space investment)\div""initial\space investment"

"1.24375 \times initial \space investment=195000+initial\space investment"

"1.24375\times initial\space investment-initial\space investment=195000"

"initial\space investment (1.24374 -1)=195000/0.24373=800065.6"

so option  [2] R800 000 appears to be the correct answer as it is close to the calculated value.